# The neutrino force at all length scales

### v1

@@ A.2 The penguin diagrams
×$\left.\quad\times\bar{v}(p_{2})\gamma^{\nu}(c_{V}^{\chi_{2}}-c_{A}^{\chi_{2}}\gamma_{5})v(p^{\prime}_{2})G^{Z}_{\mu\nu}(q)G^{W}_{\rho\sigma}(k)\right],$
$\equiv-\frac{g^{4}}{128c_{W}^{2}}\overline{u}(p^{\prime}_{1})\Gamma^{\mu}u(p_{1})\bar{v}(p_{2})\gamma^{\nu}(c_{V}^{\chi_{2}}-c_{A}^{\chi_{2}}\gamma^{5})v(p^{\prime}_{2})G^{Z}_{\mu\nu}(q)\;.$ (111)
@@
where ${\rm Li}_{2}$ is the polylogarithm function. Although $F_{1}$ is divergent and gauge dependent, it should be noticed that the discontinuity part is finite and the gauge parameter drops out:
@@
Using the dispersion technique in Eq. (15), we obtain:
$V_{0}^{\rm PG}(r)=-\frac{g^{4}}{2048\pi^{3}c_{W}^{2}r}\int_{0}^{\infty}{\rm d}t\,e^{-\sqrt{t}\,r}\,\frac{t\left(2m_{W}^{2}+3t\right)-2\left(m_{W}^{2}+t\right)^{2}\log\left(1+\frac{t}{m_{W}^{2}}\right)}{\left(t-4m_{1}^{2}\right)^{2}\left(t-m_{Z}^{2}\right)}\;.$ (120)
To get the result for the other penguin diagram in Fig. 2(c) with the vertex correction at $\overline{\chi_{2}}$, one only needs to replace $m_{1}$ with $m_{2}$ in the above equation.
At long distances $r\gg m_{Z}^{-1}$, the integral in Eq. (120) can be worked out analytically:
@@ A.3 The box diagram
$I_{0}(t)=$ $\frac{-ig^{4}}{18432m_{W}^{4}\pi^{2}t^{2}}\Bigg{\{}16t^{3}-6t^{2}\left(3m_{W}^{2}\xi_{W}-9m_{W}^{2}+t\right)\left[\Delta_{\rm E}+\log\left(\frac{\mu^{2}}{m_{W}^{2}}\right)\right]$
$+84m_{W}^{2}\xi_{W}t^{2}+6t^{2}\left(6m_{W}^{2}\xi_{W}+t\right)\left[\Delta_{\rm E}+\log\left(\frac{\mu^{2}}{m_{W}^{2}\xi_{W}}\right)\right]$
$+6t^{2}\left(4m_{W}^{2}\xi_{W}+t\right)\Lambda\left(-t,m_{W}\sqrt{\xi_{W}},m_{W}\sqrt{\xi_{W}}\right)$
$-6t\left[m_{W}^{4}+m_{W}^{2}\xi_{W}\left(m_{W}^{2}\xi_{W}-2m_{W}^{2}+2t\right)-10m_{W}^{2}t+t^{2}\right]\Lambda\left(-t,m_{W},m_{W}\sqrt{\xi_{W}}\right)$
$-2t\left[3m_{W}^{4}+3m_{W}^{2}\xi_{W}\left(m_{W}^{2}\xi_{W}-2m_{W}^{2}+7t\right)-51m_{W}^{2}t+8t^{2}\right]$
$+3\log\xi_{W}\left[m_{W}^{2}\left(-\xi_{W}+1\right)+t\right]$
×$\times\left[m_{W}^{4}+m_{W}^{2}\xi_{W}\left(m_{W}^{2}\xi_{W}-2m_{W}^{2}+4t\right)-10m_{W}^{2}t+t^{2}\right]\Bigg{\}}\;,$ (124)
@@
comes from the branch cut of Passarino-Veltman $B_{0}$ function in the $t$-plane and $\lambda(a,b,c)\equiv a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$ is the Källén kinematic function. Even though $I_{0}$ is gauge dependent and divergent, its discontinuity is not:
${\rm Disc}[I_{0}(t)]=-\frac{g^{4}}{64\pi}\frac{m_{W}^{2}t\left(t+2m_{W}^{2}\right)+\left(t^{2}-2m_{W}^{4}\right)\left(t+m_{W}^{2}\right)\log\left(1+\frac{t}{m_{W}^{2}}\right)}{t^{2}\left(t+2m_{W}^{2}\right)^{2}}\;.$ (125)

### v2

@@ A.2 The penguin diagrams
$\left.\quad\times\bar{v}(p_{2})\gamma^{\nu}(c_{V}^{\chi_{2}}-c_{A}^{\chi_{2}}\gamma_{5})v(p^{\prime}_{2})G^{Z}_{\mu\nu}(q)G^{W}_{\rho\sigma}(k)\right]$
$\equiv-\frac{g^{4}}{128c_{W}^{2}}\overline{u}(p^{\prime}_{1})\Gamma^{\mu}u(p_{1})\bar{v}(p_{2})\gamma^{\nu}(c_{V}^{\chi_{2}}-c_{A}^{\chi_{2}}\gamma^{5})v(p^{\prime}_{2})G^{Z}_{\mu\nu}(q)\;,$ (112)
@@
where ${\rm Li}_{2}$ is the dilogarithm function defined as
${\rm Li}_{2}(z)\equiv-\int_{0}^{z}{\rm d}u\,\frac{\log\left(1-u\right)}{u}\;.$
Although $F_{1}$ is divergent and gauge dependent, it should be noticed that the discontinuity part is finite and the gauge parameter drops out. In fact, the terms in Eq. (119) that contribute to the nonzero discontinuity are found to be (with $t>0$):
${\rm Disc}\left[\log\left(-\frac{m_{W}^{2}}{t}\right)\right]$ $\equiv\log\left(-\frac{m_{W}^{2}}{t+i\epsilon}\right)-\log\left(-\frac{m_{W}^{2}}{t-i\epsilon}\right)=2\pi i\;,$ (120)
${\rm Disc}\left[{\rm Li}_{2}\left(1+\frac{m_{W}^{2}}{t}\right)\right]$ $\equiv{\rm Li}_{2}\left(1+\frac{m_{W}^{2}}{t+i\epsilon}\right)-{\rm Li}_{2}\left(1+\frac{m_{W}^{2}}{t-i\epsilon}\right)=-2\pi i\log\left(1+\frac{m_{W}^{2}}{t}\right),$ (121)
${\rm Disc}\left[\log^{2}\left(-\frac{m_{W}^{2}}{t}\right)\right]$ $\equiv\log^{2}\left(-\frac{m_{W}^{2}}{t+i\epsilon}\right)-\log^{2}\left(-\frac{m_{W}^{2}}{t-i\epsilon}\right)=4\pi i\log\left(\frac{m_{W}^{2}}{t}\right).$ (122)
Combining Eqs. (120)-(122) together with (118) and (119), we obtain
@@
To get the result for the other penguin diagram in Fig. 2(c) with the vertex correction at $\overline{\chi_{2}}$, one only needs to replace $m_{1}$ with $m_{2}$ in the above calculation. At the leading order of $m_{i}^{2}/m_{Z}^{2}$ (for $i=1,2$), using the dispersion technique in Eq. (15), we obtain:
$V_{0}^{\rm PG}(r)=-\frac{g^{4}}{2048\pi^{3}c_{W}^{2}r}\int_{0}^{\infty}{\rm d}t\,e^{-\sqrt{t}\,r}\,\frac{t\left(2m_{W}^{2}+3t\right)-2\left(m_{W}^{2}+t\right)^{2}\log\left(1+\frac{t}{m_{W}^{2}}\right)}{t^{2}\left(t-m_{Z}^{2}\right)}\;.$ (124)
Note that the above integral contains a pole at $t=m_{Z}^{2}$, which comes from the $Z$ propagator in the penguin diagram. Such a divergence is superficial and can be removed by introducing the $i\epsilon$ prescription (see Appendix C for details).
At long distances $r\gg m_{Z}^{-1}$, the result can be worked out analytically:
@@ A.3 The box diagram
$I_{0}(t)=$ $-\frac{ig^{4}}{3072\pi^{2}\,t^{2}m_{W}^{4}\left(2m_{W}^{2}+t\right)^{2}}\Bigg\{-24m_{W}^{6}t\left(2m_{W}^{2}+t\right)\,\log\left(-\frac{m_{W}^{2}}{t}\right)$
$+24m_{W}^{4}t\left(m_{W}^{2}+t\right)\left(t^{2}-2m_{W}^{4}\right)\,C_{0}\left(0,0,t;0,m_{W},0\right)$
$+24m_{W}^{6}t\left(2m_{W}^{4}+6m_{W}^{2}t+3t^{2}\right)\,C_{0}\left(0,0,-t;m_{W},0,m_{W}\right)$
$-2t\left(2m_{W}^{2}+t\right)^{2}\left[t^{2}+2m_{W}^{2}t\left(\xi_{W}-5\right)+m_{W}^{4}\left(\xi_{W}-1\right)^{2}\right]\Lambda\left(-t,m_{W},m_{W}\sqrt{\xi_{W}}\right)$
$+t^{2}\left(2m_{W}^{2}+t\right)^{2}\left(4m_{W}^{2}\xi_{W}+t\right)\,\Lambda\left(-t,m_{W}\sqrt{\xi_{W}},m_{W}\sqrt{\xi_{W}}\right)$
$+t\left(2m_{W}^{2}+t\right)\left(-24m_{W}^{6}-64m_{W}^{4}t-18m_{W}^{2}t^{2}+t^{3}\right)\,\Lambda\left(-t,m_{W},m_{W}\right)$
$-m_{W}^{2}\left(2m_{W}^{2}+t\right)^{2}\log(\xi_{W})\left[3t^{2}\left(\xi_{W}+3\right)+3m_{W}^{2}t\left(\xi_{W}^{2}-4\xi_{W}+3\right)+m_{W}^{4}\left(\xi_{W}-1\right)^{3}\right]$
$-2tm_{W}^{4}\left(2m_{W}^{2}+t\right)^{2}\left(\xi_{W}^{2}-2\xi_{W}+13\right)\Bigg\}\;,$ (128)
@@
comes from the branch cut of Passarino-Veltman $B_{0}$ function in the $t$-plane and $\lambda(a,b,c)\equiv a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$ is the Källén kinematic function. The $C_{0}$ in Eq. (128) are the three-point Passarino-Veltman functions, and their explicit expressions are given by
$C_{0}\left(0,0,t;0,m_{W},0\right)=\frac{1}{t}\left[\frac{\pi^{2}}{6}+\frac{1}{2}\log^{2}\left(-\frac{m_{W}^{2}}{t}\right)+{\rm Li}_{2}\left(1+\frac{m_{W}^{2}}{t}\right)\right],$ (129)
$C_{0}\left(0,0,-t;m_{W},0,m_{W}\right)$
$=\frac{1}{t}\left[\frac{\pi^{2}}{6}+\frac{1}{2}\log^{2}\left(\frac{t-\sqrt{t\left(t+4m_{W}^{2}\right)}}{2t}\right)-\frac{1}{2}\log^{2}\left(\frac{t+2m_{W}^{2}+\sqrt{t\left(t+4m_{W}^{2}\right)}}{-t+\sqrt{t\left(t+4m_{W}^{2}\right)}}\right)\right$.
$\left.\quad\quad-{\rm Li}_{2}\left(1+\frac{t}{m_{W}^{2}}\right)-{\rm Li}_{2}\left(\frac{2\left(t+m_{W}^{2}\right)}{t-\sqrt{t\left(t+4m_{W}^{2}\right)}}\right)+{\rm Li}_{2}\left(\frac{2m_{W}^{2}}{-t+\sqrt{t\left(t+4m_{W}^{2}\right)}}\right)\right$.
$\left.\quad\quad+{\rm Li}_{2}\left(\frac{2\left(t+m_{W}^{2}\right)}{t+2m_{W}^{2}-\sqrt{t\left(t+4m_{W}^{2}\right)}}\right)-{\rm Li}_{2}\left(\frac{2m_{W}^{2}}{t+2m_{W}^{2}+\sqrt{t\left(t+4m_{W}^{2}\right)}}\right)\right].$ (130)
Then we calculate the discontinuity of $I_{0}$. First, we note that $\Lambda(-t,m_{0},m_{1})$ has no branch cut for positive $t$, so it does not contribute to the discontinuity for positive $t$. Then only the first three lines in Eq. (128) can contribute to the discontinuity. The contribution in the first line from the logarithm has been calculated in Eq. (120). For the second line in Eq. (128), using Eqs. (121) and (122) we obtain
${\rm Disc}\left[C_{0}\left(0,0,t;0,m_{W},0\right)\right]$ $\equiv C_{0}\left(0,0,t+i\epsilon;0,m_{W},0\right)-C_{0}\left(0,0,t-i\epsilon;0,m_{W},0\right)$
$=-\frac{2\pi i}{t}\log\left(1+\frac{t}{m_{W}^{2}}\right).$ (131)
Finally, for the third line in Eq. (128), an explicit calculation gives
${\rm Disc}\left[C_{0}\left(0,0,-t;m_{W},0,m_{W}\right)\right]$
$\equiv C_{0}\left(0,0,-t-i\epsilon;m_{W},0,m_{W}\right)-C_{0}\left(0,0,-t+i\epsilon;m_{W},0,m_{W}\right)=0\;.$ (132)
Therefore, we obtain the discontinuity of Eq. (128) as follows:
${\rm Disc}[I_{0}(t)]$ $=-\frac{ig^{4}}{3072\pi^{2}}\frac{-24m_{W}^{6}t\left(2m_{W}^{2}+t\right)(2\pi i)+24m_{W}^{4}\left(m_{W}^{2}+t\right)\left(t^{2}-2m_{W}^{4}\right)(-2\pi i)\log\left(1+\frac{t}{m_{W}^{2}}\right)}{m_{W}^{4}t^{2}\left(t+2m_{W}^{2}\right)^{2}}$
$=-\frac{g^{4}}{64\pi}\frac{m_{W}^{2}t\left(t+2m_{W}^{2}\right)+\left(t^{2}-2m_{W}^{4}\right)\left(t+m_{W}^{2}\right)\log\left(1+\frac{t}{m_{W}^{2}}\right)}{t^{2}\left(t+2m_{W}^{2}\right)^{2}}\;.$ (133)